use Laplace transformdx+2 dx + x = 5e»? +1 And the initial conditions are, x(0)=2 d (0)=1
Answer #1

Ginen o enital value problem da +2 dn+1=5624 del x+zat n=5024 Apply Caplace transform on both sides h{}+2h {alth ( 23 = 5ikāls)} = I can see } +4 scrol+Ky Story - Contacto 5 - ACS+1) (S+2)+B(S+2)+cesto put sarf > S = 8C-1+2) =B=5 $=-2 1 5 =0(-225+5 - A_ B (5+1) Sti istif 25+5 = A(5+1) +B put sa-/ - (358) Corellicut of 5 =) (2=A) - (sti? = get t et till = 26ttstét -